OpenStudy (anonymous):

Triangle ABC has sides with lengths 14, 17, and 26. Is the triangle obtuse, acute, or a right triangle? Answer A. Acute B. Obtuse C. Right D. Not enough information

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

@phi help

OpenStudy (anonymous):

i know its 14^2 and 17^2 and 26^ but how do i figure out its a obtuse or acute

OpenStudy (anonymous):

@sai1234

OpenStudy (anonymous):

@sai1234

OpenStudy (anonymous):

14^2+17^2=485, 26^2=676. So, the sides would have to be longer for it to be a right triangle. If you picture what that means, you can see that it is obtuse. In general, if \(a^2+b^2<c^2\), the triangle is obtuse, if \(a^2+b^2=c^2\) it is right, and if \(a^2+b^2>c^2\) it is acute.

OpenStudy (lgbasallote):

^best possible explanation

OpenStudy (phi):

A right triangle would have a longest side equal to (about) 22 

OpenStudy (phi):

If we "turn" the vertical side to the right (keep the same length), we see the longest side is shorter than 22. We have an acute triangle.